Nesmith Corporations outstanding bonds have a $1,000 par value, an 8% semiannual coupon, 20 years to maturity, and a 12% YTMq 13
Answer #1

Rate = 12% / 2 = 6%

Coupon = (0.08 * 1000) / 2 = 40

Number of periods = 20 * 2 = 40

Bond price = Coupon * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n

Bond price = 40 * [1 - 1 / (1 + 0.06)40] / 0.06 + 1000 / (1 + 0.06)40

Bond price = 40 * [1 - 0.097222] / 0.06 + 97.222188

Bond price = 40 * 15.0463 + 97.222188

Bond price = $699.07